# Deep Optical Emission-Line Images of Nine Known and Three New Galactic Supernova Remnants

### v1

@@ Abstract
Deep optical emission-line images are presented for nine known plus three new Galactic supernova remnants (SNRs), all but one having at least one angular dimension greater than one degree. Wide-field images taken in H and [O iii] 5007 reveal many new and surprising remnant structures including large remnant shock extensions and ‘blowout’ features not seen in published optical or radio data. These images represent over 12,000 individual images totaling more than 1000 hours of exposure time taken over the last two years mainly using small aperture telescopes which detected fainter nebular line emissions than published emission-line images. During the course of this imaging program, we discovered three new SNRs, namely G107.5-5.1 (the Nereides Nebula), G209.9-8.2, and G210.5+1.3, two of which have diameters . Besides offering greater structural detail on the nine already known SNRs, a key finding of this study is the importance of [O iii] emission-line imaging for mapping the complete shock emissions of Galactic SNRs.
@@ 1.1 Background and History
At present there are just over 300 confirmed Galactic supernova remnants (SNRs) cataloged (Ferrand & Safi-Harb, 2012; Green, 2019). Most are less than a degree in angular size, located within five degrees of the Galactic plane, more than 1 kpc distant, and have typical estimated ages between 10 and 10 yr (Green, 2004; Shan et al., 2018; Ranasinghe & Leahy, 2022). The vast majority are discovered through radio observations due to their characteristic synchrotron nonthermal emission associated with shocked gas leading to a power law flux density, S, with S where is the emission spectral index with typical values between 0.3 and 0.7 (Reynolds, 2011; Dubner & Giacani, 2015).
@@
Today, the identification of a Galactic SNR’s optical emission is most often done by either focused imaging searches using narrow emission-line bandpass filters (e.g., Mavromatakis et al. 2001, 2005, 2009; Boumis et al. 2002, 2009; Sezer et al. 2012; How et al. 2018; Fesen et al. 2019, 2020; Bakış et al. 2023) or using recent H surveys of the Galactic plane. Such surveys include the Virginia Tech Spectral-line Survey (VTSS; Dennison et al. 1998), the Southern H Sky Survey Atlas (SHASSA; Gaustad et al. 2001), the Wisconsin H Mapper (WHAM; Haffner et al. 2003), the AAO/UKST SuperCOSMOS H survey (Parker et al., 2005), and the Issac Newton Telescope Photometric H Survey of the Northern Galactic Plane (IPHAS; Drew et al. 2005). These surveys have served to identify optical emissions from known SNRs (e.g., Walker et al. 2001; Stupar et al. 2008; Stupar & Parker 2011) or aid in discovering new ones (Stupar et al., 2007; Stupar & Parker, 2012; Stupar et al., 2018; Sabin et al., 2013; Fesen et al., 2015, 2021)
@@
Unlike the case of steady state photoionization like that seen in H II regions, optical shock emission comes from postshock regions which have extended cool recombination zones rich in S ions. An empirical [S ii] 6716,6731 vs. H intensity ratio has proven quite useful in identifying the shocked emission of SNRs in both the Milky Way and nearby local group galaxies (Leonidaki et al., 2013; Kopsacheili et al., 2020; Long, 2017).
@@
| Name | Aperture | (degrees) | (arcsec) | Filter(s) | (FWHM Å) | Time (hr) | (Rayleighs) |
@@
| ELSMW | Nikkor 10 cm lens | 7.1 | 30 – 40 | H, [O III], [S II] | 75, 28, 50 | 0.25 | 10 – 30 |
| VTSS | Nikkor 5 cm lens | 96 | H | 17 | 2.0 | 2 | |
| SHASSA | Canon 3.3 cm lens | 48 | H | 32 | 1.7 | 0.5 – 2 | |
| WHAM | 0.6 m lens | 1 | 3600 | H | 0.25 | 0.003 | 0.15 |
| IPHAS | INT 2.5 m | 1 | H | 95 | 0.07 | 3 | |
| SHS | AAO/UKST 1.2 m | 1 | H | 80 | 3.0 | 5 | |
| Condor | cm lenses | 1 | H, [O III], [S II] | 40 | 1 – 20 | 3 – 5 | |
@@ 1.2 Wide-Field Imaging with Small Telescopes
On the other hand, small aperture telescopes with very fast optics have both much larger fields of view than large telescopes and are capable of detecting very faint emission nebulae. Since surface brightness is independent of telescope aperture, and because detected surface brightness is a function of pixel size on the sky, the number of pixels in an object, and the exposure time, long series of exposures taken with small telescopes with large FOVs and pixel scales have been shown to detect faint emission from large, extended objects as good or even better than using large telescopes (e.g., Javanmardi et al. 2016; Martinez-Delgado 2020).
@@ 1.4 Aim of this Study
Because many large angular Galactic SNRs are relatively nearby (d 3 kpc), they offer higher spatial resolution of a remnant’s interstellar shock structure. Furthermore, as a SNR ages into the later radiative phase of evolution, both their radio and optical emissions often become more filamentary due to unstable post-shock cooling leading to small scale shock features (Duin & van der Laan, 1975; Gaetz et al., 1988; Raymond et al., 2020).
@@ 2.1 Optical images
Details concerning instrument, image FOV and resolution, along with filter exposure details are given in Table 3. Our images were taken during the last two years from roughly a dozen different observing sites in Europe, Africa, New Zealand, and the USA. Over 12,000 individual emission-line filter images were obtained on these 12 objects with a total exposure time exceeding 1000 hours, or roughly 1.5 months of open shutter time (see Table 3).
@@
| Y. Sainty, N. Martino, R. Galli | Takahashi FSQ-106ED | 2.5/2.0 | [O III] | 22.0 (s) | |
@@
| G150.3+4.5 | N. Martino, Y. Sainty | Takahashi FSQ-106ED | 3.5/2.0 | [O III] | 10.0 (s) |
| Haute-Alpes, France | Takahashi FSQ-85ED | 2.7/4.6 | H+[O III] | 13.8 (s) | |
@@
| G181.1+9.5 | S. Walker | Astro-Physics 130mm | 3.4/3.2 | H | 3.9 (s) |
@@
| G209.9-8.2 | Y. Sainty, R. Galli, | Takahashi FSQ-106ED | 3.5/2.0 | H | 86.1 (s) |
| A.E. Kaeouach | Takahashi FSQ-85ED | 3.0/1.7 | [O III] | 98.7 (s) | |
@@
| B. Falls, C. Morgan | Takahashi FSQ-106ED | 2.5/2.0 | H | 35.8 (s) | |
@@
| PW Delta Rho 350 | 2.0/0.7 | [O III] | 16.6 (s) | ||
| G210.5+1.3 | M. Dreschsler, M. Blauensteiner | Lacerta 25cm | 1.2/1.1 | H | 18.0 (s) |
@@ 2.2 Optical Spectra
In a few cases, low-dispersion optical spectra of filaments of these remnants were obtained with the MDM 2.4m Hiltner telescope using Ohio State Multi-Object Spectrograph (OSMOS; Martini et al. 2011). Employing a blue VPH grism (R = 1600) and a 1.2 arcsec wide slit, exposures of s were taken covering 4000–6900 Å with a spectral resolution of 1.2 Å pixel and a FWHM = 3.5 Å. Spectra were extracted from regions clean of appreciable emission along each of the long slits.
@@ 3.1 Known Galactic SNRs
This remnant was discovered as a diffuse, soft X-ray ROSAT source by Seward et al. (1995) who reported coincident optical emission visible on the broadband red Palomar Sky Survey images along the south central boundary of the X-ray emission. Follow-up MDM 1.3 m H images revealed a complex of thin filaments extending some 20 in length along the remnant’s southern boundary.
Spectra of one optical filament confirmed its shock nature through strong [S ii] 6716,6731 ([S ii]/H = 0.84) and [O i] 6300,6364 emissions. The lack of [O iii] 4959,5007 line emission for this one filament was seen as suggesting a shock velocity 60 km s.
@@
Better views of the [O iii] filaments are shown in Figure 2 where we present close-up views of the main group of [O iii] emission filaments. Many individual filaments are visible along with extended diffuse emission444The orientation of the original top image led to unusual cropping.. The bottom panel shows a smaller FOV but higher resolution image, highlighting that the filamentary emissions are sharpest along their southern edges, marking the start of a gradual fainting of diffuse emission. This emission morphology is similar to that seen in a few other SNRs and is suggestive of an extended sheet of shocked gas viewed edge-on along the line of sight. Low dispersion spectra indicate a I([O iii] 5007)/I(H) 7 suggesting shocks in excess of 100 km s (Raymond et al., 1983, 2020).
@@
There are no published radio emission maps of G13.3-1.3 to inform us concerning possible radio-optical correlations. We therefore examined published 100 m Effelsberg surveys at 11 cm (resolution = 4.3) and 21 cm (resolution = 9.4) (Reich et al., 1990a, b) and removed the dominating large-scale Galactic emission by ‘unsharp masking’ (Sofue & Reich, 1979) from the survey maps to separate the emission from G13.3-1.3. In Figure 3, we show our [O iii] image of G13.3-1.3 (top panel) along with the 11 cm and 21 cm radio maps colored for flux in the bottom two panels.
@@
Guided by VTSS H images, an independent imaging and spectral investigation by Fesen et al. (2015) found that H filaments centered around RA = 21:24:00, Dec = +19:23 (J2000) comprised a coherent H and near-UV emission structure with coincident faint X-ray emission along its northern limb. Follow-up, low and high resolution spectra suggest a fairly nearby and old SNR (d = 1 kpc; age 90,000 yr) with modest shock speeds ( km s), the presence of [O iii] bright filaments in some regions, and an unusually thin postshock zone where preshock neutral atoms are rapidly excited and ionized (Raymond et al., 2020).
@@
The only deep optical imaging of the G82.2+5.3/W63 remnant published to date are those taken by Mavromatakis et al. (2004). They obtained flux calibrated wide images ( pixel) of the remnant in several emission lines including H + [N ii], [S ii], [O iii], and [O ii]. The W63 region of the Galactic plane in Cygnus is quite complex in terms of H emission features and the [O iii] images of Mavromatakis et al. (2004) provided the first clear picture of the G82.2+5.3 remnant, appearing markedly different to images taken in the low ionization lines of [O ii] and [S ii]. The [O iii] images revealed filamentary emission along the remnant’s west and east limbs defining the opposite sides of an ellipsoidal shell similar with the morphology seen in the remnant’s bright radio emission.
@@
The correlation between the remnant’s outlying [O iii] features with its radio emission is shown in Figure 11 where we show a positive version of our [O iii] image alongside an unpublished Effelsberg 4.8 GHz radio map (FHWM = 2.4) of the remnant. Both the [O iii] and radio images show emission off to the west/southwest of W63’s bright shell. Two bright radio features/blob correspond to seemingly unremarkable [O iii] emission clouds. In addition, we note that faint [O iii] emission located off to the shell’s north is weakly detected in the radio.
@@
A clearer picture of HB 21’s associated optical emission was made by Mavromatakis et al. (2007) who obtained wide H and [S ii] images of the remnant. They found good correlation of [S ii] strong filaments with 4850 MHz radio emission along the remnant’s eastern area. Overall, the remnant’s emission was relatively faint and patchy in the central and west areas of the remnant. These authors also obtained spectra of a half dozen regions which showed weak [O iii] emission suggesting relatively low shock velocities ( 100 km s). This conclusion was viewed in-line with the interferometric observations by Lozinskaia (1980) who observed velocities only as high as 60 to 80 km s.
@@
The remnant’s radio emission shell is incomplete along its north and northwest and more recent radio studies of the remnant (Pineault et al., 1997; Sun et al., 2011) have concluded that this gap is due to the remnant expanding into a much lower density ISM along these sections of its northern limb. X-ray data point to a center-filled SNR whose spectrum suggests the presence of a pulsar, a synchrotron nebula, plus a thermal component associated with the remnant’s expanding ISM shocks (Slane et al., 1997). The presence of a pulsar inside CTA 1 was subsequently confirmed by the Fermi Gamma-Ray Space Telescope which found a pulsar located near the center of the compact synchrotron nebula in CTA 1 (Abdo et al., 2008; Li et al., 2016). The pulsar has a period of 316.86 ms and a characteristic age of 10 yr roughly comparable to that estimated for the SNR.
@@
The fact this shell is located along the western boundary of the G150.3+4.5 remnant and exhibits similar [O iii] emission line strengths, suggests the two objects might lie at a similar distance. But this appears unlikely. CI Cam’s Gaia DR3 parallax is 0.210 mas indicating a statistical corrected distance 4.10 kpc, a value consistent with estimated based on other data (Aret et al., 2020). However, if the remnant’s distance is also kpc, its angular diameter of would indicate unrealistic physical dimensions around pc, thereby ranking it as the largest (and oldest) SNR known despite displaying an abundance of [O iii] bright filaments indicating a majority of its shocks have velocities well above 100 km s. Instead, we suggest CI Cam is unrelated to G150.3+4.5 with the SNR lying at a distance much closer in the range of kpc (Devin et al., 2020)). Consequently, the location of CI Cam so close to the western limb of G150.3+4.5 is likely an interesting but chance coincidence.
@@
Kothes et al. (2017) noted that G181.1+9.5 was the faintest SNR detected up to that time in terms of its radio emission with a surface brightness of W Hz sr at 1 GHz which put it some three times fainter than the previously faintest known SNR in 2017.
@@
This remnant was subsequently more thoroughly studied through data taken as part of Australian Square Kilometre Array Pathfinder (ASKAP) Evolutionary Map of the Universe (EMU) survey (Filipović et al., 2023). A multifrequency analysis radio data of the remnant showed it to have a thin shell morphology with a typical SNR spectral index of = and estimated its distance to be 1.3 kpc which would imply a physical diameter of around 40 pc.
@@
Instead of a small 2.5 arc minute isolated H emission cloud in the north and a tiny streak of H along the remnant’s southern boundary, G288.8-6.3’s optical is both relatively bright and quite extensive when imaged in [O iii]. We reach this conclusion even though we did not obtain a pure H image. Instead, we employed a duel bandpass filter that is sensitive to both H and [O iii] line emissions.
@@
We show a color composite of our images in Figure 26 which emphasizes the remnant’s dominated [O iii] emission shown here as blue emission. We note that we obtained the [O iii] image first and then an exploratory duel emission-line filter which suggested that a pure H image would not show many additional features. We therefore did not pursue obtaining a separate set of deep H images. As seen in other SNRs in our survey, the lack of [O iii] images of this remnant prevented previous researchers from realizing the remnant’s fuller optical emission structure.
@@
Interesting, the [O iii] image shows two ‘blowout’ features, a small one along the remnant’s northeastern tip, and a much larger but less well defined along its western and southwestern limbs. These blowout features are better seen in the enlargement presented in Figure 29. We note that neither blowout features are seen or hinted at in the radio data. While both blowouts are seen to some extent in both H and [O iii], they are best seen in [O iii] emission.

### v2

@@ Abstract
Deep optical emission-line images are presented for nine known plus three new Galactic supernova remnants (SNRs), all but one having at least one angular dimension greater than one degree. Wide-field images taken in H and [O III] 5007 reveal many new and surprising remnant structures including large remnant shock extensions and ‘breakout’ features not seen in published optical or radio data. These images represent over 12,000 individual images totaling more than 1000 hours of exposure time taken over the last two years mainly using small aperture telescopes which detected fainter nebular line emissions than published emission-line images. During the course of this imaging program, we discovered three new SNRs, namely G107.5-5.1 (the Nereides Nebula), G209.9-8.2, and G210.5+1.3, two of which have diameters . Besides offering greater structural detail on the nine already known SNRs, a key finding of this study is the importance of [O III] emission-line imaging for mapping the complete shock emissions of Galactic SNRs.
@@ 1.1 Background and History
At present there are just over 300 confirmed Galactic supernova remnants (SNRs) cataloged (Ferrand & Safi-Harb, 2012; Green, 2019). Most are less than a degree in angular size, located within five degrees of the Galactic plane, more than 1 kpc distant, and have typical estimated ages between 104 and 105 yr (Green, 2004; Shan et al., 2018; Ranasinghe & Leahy, 2022). The vast majority are discovered through radio observations due to their characteristic synchrotron nonthermal emission associated with shocked gas leading to a power law flux density, S, with S where is the emission spectral index with typical values between 0.3 and 0.7 (Reynolds, 2011; Dubner & Giacani, 2015).
@@
Today, the identification of a Galactic SNR’s optical emission is most often done by either focused imaging searches using narrow emission-line bandpass filters (e.g., Mavromatakis et al. 2001, 2005, 2009; Boumis et al. 2002, 2009; Sezer et al. 2012; How et al. 2018; Fesen et al. 2019, 2020; Bakış et al. 2023) or using recent H surveys of the Galactic plane.
Such surveys include the Marseille H Survey (MHS; Le Coarer et al. 1992; Georgelin et al. 1994, 1996, 2000; Russeil et al. 1998), the Virginia Tech Spectral-line Survey (VTSS; Dennison et al. 1998), the Southern H Sky Survey Atlas (SHASSA; Gaustad et al. 2001), the Wisconsin H Mapper (WHAM; Haffner et al. 2003), the AAO/UKST SuperCOSMOS H survey (Parker et al., 2005), and the Issac Newton Telescope Photometric H Survey of the Northern Galactic Plane (IPHAS; Drew et al. 2005). These surveys have served to identify optical emissions from known SNRs (e.g., Walker et al. 2001; Stupar et al. 2008; Stupar & Parker 2011) or aid in discovering new ones (Stupar et al., 2007; Stupar & Parker, 2012; Stupar et al., 2018; Sabin et al., 2013; Fesen et al., 2015, 2021)
@@
Unlike the case of steady state photoionization like that seen in H II regions, optical shock emission comes from postshock regions which have extended cool recombination zones rich in S+ ions. An empirical [S II] 6716,6731 vs. H intensity ratio has proven quite useful in identifying the shocked emission of SNRs in both the Milky Way and nearby local group galaxies (Leonidaki et al., 2013; Kopsacheili et al., 2020; Long, 2017).
@@
| Name | Aperture | (degrees) | (arcsec) | Filter(s) | (FWHM Å) | Time (hr) | (Rayleighs⋆) |
@@
| ELSMW1 | Nikkor 10 cm lens | 7.1 | 30 – 40 | H, [O III], [S II] | 75, 28, 50 | 0.25 | 10 – 30 |
| VTSS2 | Nikkor 5 cm lens | 96 | H | 17 | 2.0 | 2 | |
| SHASSA3 | Canon 3.3 cm lens | 48 | H | 32 | 1.7 | 0.5 – 2 | |
| WHAM4 | 0.6 m lens | 1 | 3600 | H | 0.25 | 0.003 | 0.15 |
| IPHAS5 | INT 2.5 m | 1 | H | 95 | 0.07 | 3 | |
| SHS6 | AAO/UKST 1.2 m | 1 | H | 80 | 3.0 | 5 | |
| Condor7 | cm lenses | 1 | H, [O III], [S II] | 40 | 1 – 20 | 3 – 5 | |
@@ 1.2 Wide-Field Imaging with Small Telescopes
On the other hand, small aperture telescopes with fast optics have both much larger fields of view than large telescopes and are capable of detecting very faint emission nebulae. Since surface brightness is independent of telescope aperture, and because detected surface brightness is a function of pixel size on the sky, the number of pixels in an object, and the exposure time, long series of exposures taken with small telescopes with large FOVs and pixel scales have been shown to detect faint emission from large, extended objects as good or even better than using large telescopes (e.g., Javanmardi et al. 2016; Martinez-Delgado 2020).
@@ 1.4 Aim of this Study
Because many large angular Galactic SNRs are relatively nearby (d 3 kpc), they offer higher spatial resolution of a remnant’s interstellar shock structure than more distant objects. Furthermore, as a SNR ages into the later radiative phase of evolution, both their radio and optical emissions often become more filamentary due to unstable post-shock cooling leading to small scale shock features (Duin & van der Laan, 1975; Gaetz et al., 1988; Raymond et al., 2020).
@@ 2.1 Optical images
Details concerning instrument, image FOV and resolution, along with filter exposure details are given in Table 3. Our images were taken during the last two years from ten different observing sites in Europe, Africa, New Zealand, and the USA. Over 12,000 individual emission-line filter images were obtained on these 12 objects with a total exposure time exceeding 1000 hours, or roughly 1.5 months of open shutter time (see Table 3).
@@
| Y. Sainty, N. Martino, R. Galli | Takahashi FSQ-106ED | 2.5/2.0′′ | [O III] | 22.0 (s) | |
@@
| G150.3+4.5 | N. Martino, Y. Sainty | Takahashi FSQ-106ED | 3.5/2.0′′ | [O III] | 10.0 (s) |
| Haute-Alpes, France | Takahashi FSQ-85ED | 2.7/4.6′′ | H+[O III] | 13.8 (s) | |
@@
| G181.1+9.5 | S. Walker | Astro-Physics 130mm | 3.4/3.2′′ | H | 3.9 (s) |
@@
| G209.9-8.2 | Y. Sainty, R. Galli, | Takahashi FSQ-106ED | 3.5/2.0′′ | H | 86.1 (s) |
| A.E. Kaeouach | Takahashi FSQ-85ED | 3.0/1.7′′ | [O III] | 98.7 (s) | |
@@
| B. Falls, C. Morgan | Takahashi FSQ-106ED | 2.5/2.0′′ | H | 35.8 (s) | |
@@
| PW Delta Rho 350 | 2.0/0.7′′ | [O III] | 16.6 (s) | ||
| G210.5+1.3 | M. Dreschsler, M. Blauensteiner | Lacerta 25cm | 1.2/1.1′′ | H | 18.0 (s) |
@@ 2.2 Optical Spectra
In a few cases, low-dispersion optical spectra of filaments of these remnants were obtained with the MDM 2.4m Hiltner telescope using Ohio State Multi-Object Spectrograph (OSMOS; Martini et al. 2011). Employing a blue VPH grism (R = 1600) and a 1.2 arcsec wide slit, exposures of s were taken covering 4000–6900 Å with a spectral resolution of 1.2 Å pixel-1 and a FWHM = 3.5 Å. Spectra were extracted from regions clean of appreciable emission along each of the long slits.
@@ 3.1 Known Galactic SNRs
This remnant was discovered as a diffuse, soft X-ray ROSAT source by Seward et al. (1995) who reported coincident optical emission visible on the broadband red Palomar Sky Survey images along the south central boundary of the X-ray emission. Follow-up MDM 1.3 m H images revealed a complex of thin filaments extending some 20′ in length along the remnant’s southern boundary.
Spectra of one optical filament confirmed its shock nature through strong [S II] 6716,6731 ([S II]/H = 0.84) and [O I] 6300,6364 emissions. The lack of [O III] 4959,5007 line emission for this one filament was seen as suggesting a shock velocity 60 km s-1.
@@
Better views of the [O III] filaments are shown in Figure 2 where we present close-up views of the main group of [O III] emission filaments. Many individual filaments are visible along with extended diffuse emission444The orientation of the original top image led to unusual cropping.. The bottom panel shows a smaller FOV but higher resolution image, highlighting that the filamentary emissions are sharpest along their southern edges, marking the start of a gradual fainting of diffuse emission. This emission morphology is similar to that seen in a few other SNRs and is suggestive of an extended sheet of shocked gas viewed edge-on along the line of sight. Low dispersion spectra indicate a I([O III] 5007)/I(H) 7 suggesting shocks in excess of 100 km s-1 (Raymond et al., 1983, 2020).
@@
There are no published radio emission maps of G13.3-1.3 to inform us concerning possible radio-optical correlations. We therefore examined published 100 m Effelsberg surveys at 11 cm (resolution = 4.3′) and 21 cm (resolution = 9.4′) (Reich et al., 1990a, b) and removed the dominating large-scale Galactic emission by ‘unsharp masking’ (Sofue & Reich, 1979) from the survey maps to separate the emission from G13.3-1.3. In Figure 3, we show our [O III] image of G13.3-1.3 (top panel) along with the 11 cm and 21 cm radio maps colored for flux in the bottom two panels.
@@
Guided by VTSS H images, an independent imaging and spectral investigation by Fesen et al. (2015) found that H filaments centered around RA = 21:24:00, Dec = +19:23 (J2000) comprised a coherent H and near-UV emission structure with coincident faint X-ray emission along its northern limb. Follow-up, low and high resolution spectra suggest a fairly nearby and old SNR (d = 1 kpc; age 90,000 yr) with modest shock speeds ( km s-1), the presence of [O III] bright filaments in some regions, and an unusually thin postshock zone where preshock neutral atoms are rapidly excited and ionized (Raymond et al., 2020).
@@
The only deep optical imaging of the G82.2+5.3/W63 remnant published to date are those taken by Mavromatakis et al. (2004). They obtained flux calibrated wide images ( pixel-1) of the remnant in several emission lines including H + [N II], [S II], [O III], and [O II]. The W63 region of the Galactic plane in Cygnus is quite complex in terms of H emission features and the [O III] images of Mavromatakis et al. (2004) provided the first clear picture of the G82.2+5.3 remnant, appearing markedly different to images taken in the low ionization lines of [O II] and [S II]. The [O III] images revealed filamentary emission along the remnant’s west and east limbs defining the opposite sides of an ellipsoidal shell similar with the morphology seen in the remnant’s bright radio emission.
@@
The correlation between the remnant’s outlying [O III] features with its radio emission is shown in Figure 11 where we show a positive version of our [O III] image alongside an unpublished Effelsberg 4.8 GHz radio map (FHWM = 2.4′) of the remnant. Both the [O III] and radio images show emission off to the west/southwest of W63’s bright shell. Two bright radio features/blob correspond to seemingly unremarkable [O III] emission clouds. In addition, we note that faint [O III] emission located off to the shell’s north is weakly detected in the radio.
@@
A clearer picture of HB 21’s associated optical emission was made by Mavromatakis et al. (2007) who obtained wide H and [S II] images of the remnant. They found good correlation of [S II] strong filaments with 4850 MHz radio emission along the remnant’s eastern area. Overall, the remnant’s emission was relatively faint and patchy in the central and west areas of the remnant. These authors also obtained spectra of a half dozen regions which showed weak [O III] emission suggesting relatively low shock velocities ( 100 km s-1). This conclusion was viewed in-line with the interferometric observations by Lozinskaia (1980) who observed velocities only as high as 60 to 80 km s-1.
@@
The remnant’s radio emission shell is incomplete along its north and northwest and more recent radio studies of the remnant (Pineault et al., 1997; Sun et al., 2011) have concluded that this gap is due to the remnant expanding into a much lower density ISM along these sections of its northern limb. X-ray data point to a center-filled SNR whose spectrum suggests the presence of a pulsar, a synchrotron nebula, plus a thermal component associated with the remnant’s expanding ISM shocks (Slane et al., 1997). The presence of a pulsar inside CTA 1 was subsequently confirmed by the Fermi Gamma-Ray Space Telescope which found a pulsar located near the center of the compact synchrotron nebula in CTA 1 (Abdo et al., 2008; Li et al., 2016). The pulsar has a period of 316.86 ms and a characteristic age of 104 yr roughly comparable to that estimated for the SNR.
@@
The fact this shell is located along the western boundary of the G150.3+4.5 remnant and exhibits similar [O III] emission line strengths, suggests the two objects might lie at a similar distance. But this appears unlikely. CI Cam’s Gaia DR3 parallax is 0.210 mas indicating a statistical corrected distance 4.10 kpc, a value consistent with estimated based on other data (Aret et al., 2020). However, if the remnant’s distance is also kpc, its angular diameter of would indicate unrealistic physical dimensions around pc, thereby ranking it as the largest (and oldest) SNR known despite displaying an abundance of [O III] bright filaments indicating a majority of its shocks have velocities well above 100 km s-1. Instead, we suggest CI Cam is unrelated to G150.3+4.5 with the SNR lying at a distance much closer than that of CI Cam (Devin et al., 2020; Feng et al., 2024)). Consequently, the location of CI Cam so close to the western limb of G150.3+4.5 is likely to be a chance coincidence.
@@
Kothes et al. (2017) noted that G181.1+9.5 was the faintest SNR detected up to that time in terms of its radio emission with a surface brightness of W-2 Hz-1 sr-1 at 1 GHz which put it some three times fainter than the previously faintest known SNR in 2017.
@@
This remnant was subsequently more thoroughly studied through data taken as part of Australian Square Kilometre Array Pathfinder (ASKAP) Evolutionary Map of the Universe (EMU) survey (Filipović et al., 2023). A multifrequency analysis radio data of the remnant showed it to have a thin shell morphology with a typical SNR spectral index of = and estimated its distance to be 1.3 kpc which would imply a physical diameter of around 40 pc. The remnant has also been detected in -rays with the -Large Area Telescope (Burger-Scheidlin et al., 2023).
@@
Instead of a small 2.5 arc minute isolated H emission cloud in the north and a tiny streak of H along the remnant’s southern boundary, G288.8-6.3’s optical is both relatively bright and quite extensive when imaged in [O III]. We reach this conclusion even though we did not obtain a pure H image. Instead, we employed a dual bandpass filter that is sensitive to both H and [O III] line emissions.
@@
We show a color composite of our images in Figure 26 which emphasizes the remnant’s dominated [O III] emission shown here as blue emission. We note that we obtained the [O III] image first and then an exploratory dual emission-line filter which suggested that a pure H image would not show many additional features. We therefore did not pursue obtaining a separate set of deep H images. As seen in other SNRs in our survey, the lack of [O III] images of this remnant prevented previous researchers from realizing the remnant’s fuller optical emission structure.
@@
Bottom: Color composite of the SNR G321.3-3.9. Blue: [O III], red: [S II], yellow: H. North is up, east to the left.
@@
Interesting, the [O III] image shows two ‘breakout’ features, a small one along the remnant’s northeastern tip, and a much larger but less well defined along its western and southwestern limbs. These breakout features are better seen in the enlargement presented in Figure 29. We note that neither breakout features are seen or hinted at in the radio data. While both breakouts are seen to some extent in both H and [O III], they are best seen in [O III] emission.